# TopoTune: A Framework for Generalized Combinatorial Complex Neural Networks

### v4

@@ B.3.2 Weisfeiler-Leman (WL) tests on Combinatorial Complexes
    $c_{\mathcal{N}}^{t}(\sigma)$ =$=\left\{\left\{c_{\sigma^{\prime},\ell}^{t}~{}|~{}\forall\sigma^{\prime}\in\mathcal{N}(\sigma)\right\}\right\},$
@@
Next, we generalize the set-based $k$-WL test to labeled combinatorial complexes, called the $k$-CCWL test. The set-based $k$-WL test is employed in (Morris et al., 2019) where colors are defined on $k$-sets of nodes, as opposed to $k$-tuples of nodes in the standard $k$-WL test. Specifically, we denote $[\mathcal{C}]^{k}$ the set of $k$-sets formed with cells of $\mathcal{C}$. We generalize the definition of neighborhood of $k$-sets of vertices from (Morris et al., 2019) to neighborhood of $k$-sets of cells.
@@
Given a $k$-set of cells $s=\left\{\sigma_{1},\ldots,\sigma_{k}\right\}$ in $[\mathcal{C}]^{k}$, we define its neighborhood as the function $\mathcal{N}_{k}:[\mathcal{C}]^{k}\mapsto\mathcal{P}([\mathcal{C}]^{k})$ defined as:
@@
  • Initialization: Every $k$-set $s$ in $[\mathcal{C}]^{k}$ is initialized with a color that corresponds to the CC-isomorphism type of the sub-CC defined by $s=\{\sigma_{1},\ldots,\sigma_{k}\}$ induced by $\mathcal{N}|_{s}$ where $\mathcal{N}|_{s}$ is the neighborhood $\mathcal{N}$ restricted to $s$. This means that two $k$-sets $s$ and $s^{\prime}$ get the same color if and only if there is a labeled CC-isomorphism (for labeling function $\ell$) between the sub-CCs corresponding to the cells in $s$ and $s^{\prime}$, respectively.
@@
    $c_{\mathcal{N}_{k}(s),\ell}^{t}$ =$=\left\{\left\{c_{s^{\prime},\ell}^{t}~{}|~{}\forall s^{\prime}\in\mathcal{N}_{k}(s)\right\}\right\},$
@@ B.3.3 Definitions of $k$-GNNs and $k$-CCNNs
We generalize the definition of $k$-GNNs by (Morris et al., 2019) into a definition of $k$-CCNNs.
@@
Let $(\mathcal{C},\ell)$ be a labeled CC. In each $k$-CCNN layer $t$, the feature vector $h^{(t)}_{k}(s)\in\mathbb{R}^{d}$ for each $k$-set $s$ in $[\mathcal{C}]^{k}$ is updated into $h^{(t+1)}_{k}(s)$ as follows:
@@ B.3.4 Relationships between CCWL/GCWL tests and CCNNs/GCCNs
We prove relationships between the expressivity of the WL tests and the expressivity of the corresponding neural networks. We first recall results on WL tests on graphs and GNNs (Morris et al., 2019). In what follows, $(G,\ell)$ is a labeled graph, and $W^{(t)}$ denote the parameters of a GNN up to layer $t$. We encode the initial labels $\ell(v)$, for a vertex $v$, by vectors $h^{(0)}(v)\in\mathbb{R}^{1\times d}$.
@@
Theorem 1 in (Morris et al., 2019) states that, for every encoding of the graph labels $\ell(v)$ as $d$-vectors $h^{(0)}(v)$, and for every choice of parameters $W^{(t)}$, the coloring $c(t)_{\ell}$ of the WL test always refines the coloring $h(t)$ induced by the GNN parameterized by $W^{(t)}$. Theorem 2 in (Morris et al., 2019) states that there exists parameter matrices $W^{(t)}$ such that GNNs have exactly the same power as the WL test. Consequently, we say that GNNs have the same expressivity as the WL test. Similarly, Propositions 3 and 4 from (Morris et al., 2019) show that $k$-GNNs have the same expressivity as the k-WL test.
@@
Equivalence of refinements. By construction of the strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}$, $k$-sets of nodes in $\mathcal{G}_{\mathcal{N}}$ are $k$-sets of cells in $\mathcal{C}_{\mathcal{N}}$, and the neighborhoods of $k$-sets of nodes defined in (Morris et al., 2019) are the neighborhoods of $k$-sets of cells defined in Definition B.7. Thus, the refinement equation of the $k$-CCWL test is equal to the refinement equation of the $k$-WL test on $\mathcal{G}_{\mathcal{N}}$.
@@
Given the equivalence between the computations in $\mathcal{C}$ and in $\mathcal{G}_{\mathcal{N}}$ provided by Proposition B.11, we can pull the results from Theorems 1 and 2 from (Morris et al., 2019) and provide the following propositions.
@@
Propositions B.15 and B.16 are given by Proposition B.12 and Propositions 3 and 4 from Morris et al. (2019). They show that $k$-CCNNs have the same power as the $k$-CCWL.
@@ B.3.5 Proof
By Proposition B.10, GCCN have at least the same expressive power as $k$-CCNNs. By Propositions B.15-B.16, $k$-CCNNs have the same expressive power as the $k$-CCWL. By Proposition B.12, the $k$-CCWL test is equivalent to the $k$-WL test on the associated strictly augmented Hasse graph. It is known (e.g., (Grohe, 2017)) that the $k$-WL test on graph is strictly more powerful than the WL test. Thus, the $k$-WL test on the strictly augmented Hasse graph is strictly more powerful than the WL test on that same graph. By Proposition B.11, the WL test on the strictly augmented Hasse graph is equivalent to the CCWL test on the corresponding CC. By Propositions B.13-B.14, the CCWL test on the CC has the same expressive power as CCNNs.
@@
  • It is known that regular graphs of the same order are indistinguishable by the WL test (see, e.g., (Kiefer, 2020; Morris et al., 2023)).
  • Every pair of graphs with $n$ nodes are distinguishable by the $n$-WL test (Morris et al., 2023).
@@ C.1 Key Definitions
    where $\mathbf{h}_{y}$ is a 1×$F$ vector, $\Theta$ is an $F$×$F$ weight matrix, and $a_{xy}$ is a scalar. This involves a matrix-vector multiplication, contributing a complexity of $O(F^{2})$ per message.
@@ C.2 Complexity of $\omega_{\mathcal{N}}$
  • $A$: Complexity of aggregation (e.g., assuming sum/average, $O(F)$),
@@ E.2 Datasets
Table 3 provides the statistics for each dataset lifted to three topological domains: simplicial complex, cellular complex, and hypergraph. The table shows the number of 00-cells (nodes), $1$-cells (edges), and $2$-cells (faces) of each dataset after the topology lifting procedure. We recall that:
@@
Dataset Domain # 00-cell # $1$-cell # $2$-cell
@@
To address these limitations, ongoing research is focused on developing scalable lifting procedures that can extend TDL methods to broader datasets, including those with more complex structures or larger scales. For example, Bernárdez et al. (2024) propose innovative topological liftings, paving the way for more scalable and applicable datasets in TDL.
@@ E.3 Hyperparameter search
Five splits are generated for each dataset to ensure a fair evaluation of the models across domains. Each split comprises 50% training data, 25% validation data, and 25% test data. An exception is made for the ZINC dataset, where predefined splits are used (Irwin et al., 2012).
@@ Appendix G Model Training Time
CCNN (Best Model on TopoBench) 100 ± 23 132 ± 19 238 ± 89 254 ± 39 228 ± 44 75 ± 15 57 ± 4.4 128 ± 50
@@
CCNN (Best Model on TopoBench) 123 ± 57 104 ± 28 172 ± 50 183 ± 62 178 ± 86 143 ± 16 75 ± 23 114 ± 18
@@
CCNN (Best Model on TopoBench) 127 ± 48 96 ± 20 220 ± 74 128 ± 49 387 ± 105 121 ± 38 48 ± 1 177 ± 71
@@
SCCN (Yang et al., 2022)
Benchmark results (Telyatnikov et al., 2024) 11 ± 2 60 ± 18 247 ± 65 311 ± 83 102 ± 39 101 ± 41 143 ± 35
@@
CWN (Bodnar et al., 2021a)
Benchmark results (Telyatnikov et al., 2024) 11 ± 2 43 ± 5 240 ± 50 252 ± 92 54 ± 25 52 ± 5 119 ± 14
@@ H.1 Larger node-level datasets
Table 8 additionally presents the experimental results on 4 heterophilic datasets introduced in (Platonov et al., ) (Amazon Ratings, Roman Empire, Minesweeper, and Questions). These represent larger node-level classification tasks than those shown in the main Table 1, with up to 48,921 nodes and 153,540 edges in the case of the Questions graph. Except on this precise dataset, which was not considered in previous TDL literature, we compare the results against CCNNs and hypergraph models from Telyatnikov et al. (2024). We observe that overall GCCNs achieve similar performance than regular CCNNs, and they outperform them by a significant margin on Minesweeper.
@@
Table 8: Results on larger node level datasets, each experiment run with 5 seeds. We report accuracy for Amazon Ratings and Roman Empire, and AUC-ROC for Minesweeper and Questions. The values for the best CCNNs and hypergraph models are extracted from TopoBench (Telyatnikov et al., 2024).
@@ H.2 More advanced GNNs
We include a subset of experiments with the same protocol as in Table 1 using GATv2 (Brody et al., 2021) and PNA (Corso et al., 2020) in the cellular domain. Results show how the GCCNs built with these models perform consistently well across node-level and graph-level tasks on the cell domain, often $<1\sigma$ of the best (standard-GNN) GCCN as reported in Table 1, but only outperform them on MUTAG.
Table 9: Cross-domain, cross-task, cross-expansion, and cross-$\omega_{\mathcal{N}}$ comparison of GCCN architectures built with GATv2 (Brody et al., 2021) and PNA (Corso et al., 2020) with top-performing GCCNs from Table 1 and benchmarked on TopoBench (Telyatnikov et al., 2024). Best result is in bold and results within 1 standard deviation are highlighted blue. Experiments are run with 5 seeds. We report accuracy for classification tasks and MAE for regression.

### v5

@@ B.3.2 Weisfeiler-Leman (WL) tests on Combinatorial Complexes
    $c_{\mathcal{N}}^{t}(\sigma)$ $=\left\{\left\{c_{\sigma^{\prime},\ell}^{t}~|~\forall\sigma^{\prime}\in\mathcal{N}(\sigma)\right\}\right\},$
@@
Next, we generalize the set-based $k$-WL test to labeled combinatorial complexes, called the $k$-CCWL test. The set-based $k$-WL test is employed in (morris2019weisfeiler) where colors are defined on $k$-sets of nodes, as opposed to $k$-tuples of nodes in the standard $k$-WL test. Specifically, we denote $\mathcal{C}]^{k}$ the set of $k$-sets formed with cells of $\mathcal{C}$. We generalize the definition of neighborhood of $k$-sets of vertices from (morris2019weisfeiler) to neighborhood of $k$-sets of cells.
@@
Given a $k$-set of cells $s=\left\{\sigma_{1},\ldots,\sigma_{k}\right\}$ in $\mathcal{C}]^{k}$, we define its neighborhood as the function $\mathcal{N}_{k}:[\mathcal{C}]^{k}\mapsto\mathcal{P}([\mathcal{C}]^{k})$ defined as:
@@
  • Initialization: Every $k$-set $s$ in $\mathcal{C}]^{k}$ is initialized with a color that corresponds to the CC-isomorphism type of the sub-CC defined by $s=\{\sigma_{1},\ldots,\sigma_{k}\}$ induced by $\mathcal{N}|_{s}$ where $\mathcal{N}|_{s}$ is the neighborhood $\mathcal{N}$ restricted to $s$. This means that two $k$-sets $s$ and $s^{\prime}$ get the same color if and only if there is a labeled CC-isomorphism (for labeling function $\ell$) between the sub-CCs corresponding to the cells in $s$ and $s^{\prime}$, respectively.
@@
    $c_{\mathcal{N}_{k}(s),\ell}^{t}$ $=\left\{\left\{c_{s^{\prime},\ell}^{t}~|~\forall s^{\prime}\in\mathcal{N}_{k}(s)\right\}\right\},$
@@ B.3.3 Definitions of $k$-GNNs and $k$-CCNNs
We generalize the definition of $k$-GNNs by (morris2019weisfeiler) into a definition of $k$-CCNNs.
@@
Let $(\mathcal{C},\ell)$ be a labeled CC. In each $k$-CCNN layer $t$, the feature vector $h^{(t)}_{k}(s)\in\mathbb{R}^{d}$ for each $k$-set $s$ in $\mathcal{C}]^{k}$ is updated into $h^{(t+1)}_{k}(s)$ as follows:
@@ B.3.4 Relationships between CCWL/GCWL tests and CCNNs/GCCNs
We prove relationships between the expressivity of the WL tests and the expressivity of the corresponding neural networks. We first recall results on WL tests on graphs and GNNs (morris2019weisfeiler). In what follows, $(G,\ell)$ is a labeled graph, and $W^{(t)}$ denote the parameters of a GNN up to layer $t$. We encode the initial labels $\ell(v)$, for a vertex $v$, by vectors $h^{(0)}(v)\in\mathbb{R}^{1\times d}$.
@@
Theorem 1 in (morris2019weisfeiler) states that, for every encoding of the graph labels $\ell(v)$ as $d$-vectors $h^{(0)}(v)$, and for every choice of parameters $W^{(t)}$, the coloring $c(t)_{\ell}$ of the WL test always refines the coloring $h(t)$ induced by the GNN parameterized by $W^{(t)}$. Theorem 2 in (morris2019weisfeiler) states that there exists parameter matrices $W^{(t)}$ such that GNNs have exactly the same power as the WL test. Consequently, we say that GNNs have the same expressivity as the WL test. Similarly, Propositions 3 and 4 from (morris2019weisfeiler) show that $k$-GNNs have the same expressivity as the k-WL test.
@@
Equivalence of refinements. By construction of the strictly augmented Hasse graph $\mathcal{G}_{\mathcal{N}}$, $k$-sets of nodes in $\mathcal{G}_{\mathcal{N}}$ are $k$-sets of cells in $\mathcal{C}_{\mathcal{N}}$, and the neighborhoods of $k$-sets of nodes defined in (morris2019weisfeiler) are the neighborhoods of $k$-sets of cells defined in Definition B.7. Thus, the refinement equation of the $k$-CCWL test is equal to the refinement equation of the $k$-WL test on $\mathcal{G}_{\mathcal{N}}$.
@@
Given the equivalence between the computations in $\mathcal{C}$ and in $\mathcal{G}_{\mathcal{N}}$ provided by Proposition B.11, we can pull the results from Theorems 1 and 2 from (morris2019weisfeiler) and provide the following propositions.
@@
Propositions B.15 and B.16 are given by Proposition B.12 and Propositions 3 and 4 from morris2019weisfeiler. They show that $k$-CCNNs have the same power as the $k$-CCWL.
@@ B.3.5 Proof
By Proposition B.10, GCCN have at least the same expressive power as $k$-CCNNs. By Propositions B.15-B.16, $k$-CCNNs have the same expressive power as the $k$-CCWL. By Proposition B.12, the $k$-CCWL test is equivalent to the $k$-WL test on the associated strictly augmented Hasse graph. It is known (e.g., (grohe2017descriptive)) that the $k$-WL test on graph is strictly more powerful than the WL test. Thus, the $k$-WL test on the strictly augmented Hasse graph is strictly more powerful than the WL test on that same graph. By Proposition B.11, the WL test on the strictly augmented Hasse graph is equivalent to the CCWL test on the corresponding CC. By Propositions B.13-B.14, the CCWL test on the CC has the same expressive power as CCNNs.
@@
  • It is known that regular graphs of the same order are indistinguishable by the WL test (see, e.g., (kiefer2020power; morris2023weisfeiler)).
  • Every pair of graphs with $n$ nodes are distinguishable by the $n$-WL test (morris2023weisfeiler).
@@ C.1 Key Definitions
    where $\mathbf{h}_{y}$ is a 1×$F$ vector, $\Theta$ is an FF×$F$ weight matrix, and $a_{xy}$ is a scalar. This involves a matrix-vector multiplication, contributing a complexity of $O(F^{2})$ per message.
@@ C.2 Complexity of $\omega_{\mathcal{N}}$
  • AA: Complexity of aggregation (e.g., assuming sum/average, $O(F)$),
@@ E.2 Datasets
Table 3 provides the statistics for each dataset lifted to three topological domains: simplicial complex, cellular complex, and hypergraph. The table shows the number of 0-cells (nodes), $1$-cells (edges), and $2$-cells (faces) of each dataset after the topology lifting procedure. We recall that:
@@
Dataset Domain # 0-cell # $1$-cell # $2$-cell
@@
To address these limitations, ongoing research is focused on developing scalable lifting procedures that can extend TDL methods to broader datasets, including those with more complex structures or larger scales. For example, challenge propose innovative topological liftings, paving the way for more scalable and applicable datasets in TDL.
@@ E.3 Hyperparameter search
Five splits are generated for each dataset to ensure a fair evaluation of the models across domains. Each split comprises 50% training data, 25% validation data, and 25% test data. An exception is made for the ZINC dataset, where predefined splits are used (irwin2012zinc).
@@ Appendix G Model Training Time
CCNN (Best Model on TopoBench) 19.8 ± 6 53 ± 5 293 ± 95 318 ± 47 1206 ± 206 52 ± 9 63 ± 5 174 ± 50
@@
CCNN (Best Model on TopoBench) 21 ± 5 49 ± 10 263 ± 63 291 ± 81 1403 ± 605 124 ± 15 99 ± 26 155 ± 21
@@
CCNN (Best Model on TopoBench) 14 ± 2 24 ± 2 130 ± 36 137 ± 38 396 ± 95 70 ± 18 54 ± 9 138 ± 43
@@
SCCN (yang2022efficientrlsimplicial)
Benchmark results (telyatnikov2024topobench) 11 ± 2 60 ± 18 247 ± 65 311 ± 83 102 ± 39 101 ± 41 143 ± 35
@@
CWN (bodnar2021weisfeiler)
Benchmark results (telyatnikov2024topobench) 11 ± 2 43 ± 5 240 ± 50 252 ± 92 54 ± 25 52 ± 5 119 ± 14
@@ H.1 Larger node-level datasets
Table 8 additionally presents the experimental results on 4 heterophilic datasets introduced in (platonov2023critical) (Amazon Ratings, Roman Empire, Minesweeper, and Questions). These represent larger node-level classification tasks than those shown in the main Table 1, with up to 48,921 nodes and 153,540 edges in the case of the Questions graph. Except on this precise dataset, which was not considered in previous TDL literature, we compare the results against CCNNs and hypergraph models from telyatnikov2024topobench. We observe that overall GCCNs achieve similar performance than regular CCNNs, and they outperform them by a significant margin on Minesweeper.
@@
Table 8: Results on larger node level datasets, each experiment run with 5 seeds. We report accuracy for Amazon Ratings and Roman Empire, and AUC-ROC for Minesweeper and Questions. The values for the best CCNNs and hypergraph models are extracted from TopoBench (telyatnikov2024topobench).
@@ H.2 More advanced GNNs
We include a subset of experiments with the same protocol as in Table 1 using GATv2 (brody2021attentive) and PNA (corso2020pna) in the cellular domain. Results show how the GCCNs built with these models perform consistently well across node-level and graph-level tasks on the cell domain, often $1\sigma$ of the best (standard-GNN) GCCN as reported in Table 1, but only outperform them on MUTAG.
Table 9: Cross-domain, cross-task, cross-expansion, and cross-$\omega_{\mathcal{N}}$ comparison of GCCN architectures built with GATv2 (brody2021attentive) and PNA (corso2020pna) with top-performing GCCNs from Table 1 and benchmarked on TopoBench (telyatnikov2024topobench). Best result is in bold and results within 1 standard deviation are highlighted blue. Experiments are run with 5 seeds. We report accuracy for classification tasks and MAE for regression.